# Diffusion Transformer Policy

### v1

@@ 4.3 CALVIN
Comparisons. Table 2 presents the comparisons with previous methods on Calvin and the proposed methods. Without whistles and bells, the proposed Diffusion Tranformer Policy achieves the state-of-the-art results. Particularly, we only use RGB camera stream for observation. The superior demonstrates the effectiveness of Diffusion Transformer Policy. Meanwhile, the pretraining on Open X-Embodiment Datasets significantly facilitates the performance by 1.23, which demonstrates the transferability of Diffusion Transformer Policy. By contrast, the performance of diffusion action head is worse than Diffusion Transformer Policy by 0.45, though we load similar pretraining weights for diffusion head architecture. Diffusion Transformer Policy can scale across different environments, e.g., transferring the knowledge from the diverse real datasets to the CALVIN dataset.
Table 2: The comparisons with state-of-the-art approaches on Calvin Benchmark under success rate and average success length. ’S’ indicates Static RGB, ’G’ indicate Gripper RGB. ’SD’ indicate Static RGB-D, ’GD’ indicates Gripper RGB-D, ’P’ indicates Proprio, i.e., the observation arm position, ’C’ indicates camera parameters.
@@
SPIL  Zhou et al. (2024) S,G 74.2% 46.3% 27.6% 14.7% 8.0% 1.71
RoboFlamingo Li et al. (2023) S,G 82.4% 61.9% 46.6% 33.1% 23.5% 2.47
SuSIE Black et al. (2023) S 87.0% 69.0% 49.0% 38.0% 26.0% 2.69
GR-1 Wu et al. (2023) S,G,P 85.4% 71.2% 59.6% 49.7% 40.1% 3.06
3D Diffuser Ke et al. (2024) SD,GD,P,C 92.2% 78.7% 63.9% 51.2% 41.2% 3.27
diffusion head w/o pretrain S 75.5% 44.8% 25.0% 15.0% 7.5% 1.68
diffusion head S 94.3% 77.5% 62.0% 48.3% 34.0% 3.16
Ours w/o pretrain S 89.5% 63.3% 39.8% 27.3% 18.5% 2.38
Ours S 94.5% 82.5% 72.8% 61.3% 50.0% 3.61
@@ 4.4 Real Franka Arm
We finally evaluate the proposed method on the Real Franka Arm. Specifically, we train the model on Open X-Embodiment Datasets Padalkar et al. (2023), and evaluate it in our Franka Arm environment under zero-shot generalization and few-shot finetuning generalization.
@@
setup. We set up the franka on the table with a black background. Meanwhile, we use a single third-person RGB camera about 1.5 meters away from the Franka Arm. Please refer to Figure 3 for the visualized demonstration. Considering the environment of our Franka setup is different from the scenes in Open X-Embodiment Stone et al. (2023), we mainly evaluate the proposed method on out-of-the-box generation and few-shot generation. In our experiments, we evaluate each model with the same scene, and the object is placed in 9 similar positions in a 9-grid format in front of the franka arm. Meanwhile, we maintain a small variance in those positions placing the objects for evaluation.
@@
To evaluate the ability of the proposed model on new environments with a few demonstrations. We set 5 tasks, including “Pick up the green block”, “Pick up the kiwifruit”, “Pick up the banana”, “Pick up the tiny green block”, “Pick up the pink block”. Meanwhile, we collect 50 trajectories for each task in the first three tasks, while leaving the remaining two tasks (“Pick up the tiny green block” and “Pick up the pink block”) for out-of-distribution evaluation. Figure 3 presents the scenes and tasks. The image in Figure 3 is the exact image for our model. Our real-world environment is challenging since the object is small compared to the whole scene.
Figure 3: Illustration of Franka environment and corresponding objects.
@@
Finetuning details. In our experiments, we finetune the proposed method on the real Franka Arm with Lora Hu et al. (2021) and AdamW  for 10,000 steps. We set the number of timesteps as 100 for DDPM Ho et al. (2020), and batch size as 512. Meanwhile, we finetune all the networks with one observation and one step prediction.
Zero-shot and 10-shot Generalization We directly take the models pretrained on Open X-Embodiment to evaluate zero-shot and 10-shot generalization in our environments. We compare the proposed method with the discretization action head baseline Brohan et al. (2022) (RT-1) that are optimized on the same dataset mixtures, as well as the released OpenVla Kim et al. (2024) and Octo-base Team et al. (2024) models. Besides, we evaluate it on three simple tasks: “pick up the green block”, “pick up the banana”, and “pick up the kiwifruit”. Table 3 shows OpenVLA and Octo without in-domain finetuning fail to pick up the object, while the proposed method can pick the green block with a success rate of 10%. We think it is because the proposed model is able to scale better the mixture data from Open X-Embodiment. Meanwhile, we find we still achieve similar performance to OpenVLA Kim et al. (2024) (a 7B model) when finetuning it with only 10 samples on task “pick up the banana”. We provide a visualized analysis in Appendix C.1, where we find the OpenVLA and Octo fail to pick due to wrong grasp pose, while the proposed model can grasp the object with a small success rate.
Table 3: Comparision with Baseline, Octo and Openvla on Real Franka Arm (pick up tasks) under success rate. Here the baseline is the model with discretization actions, similar to RT-1 Brohan et al. (2022), that we implement and train on Open X-Embodiment. 10-shot finetuning indicates we finetune the model with only 10 samples for each task.
Method Zero-shot 10-shot
Pick Block Pick Banana Pick Kiwifruit Pick Block Pick Banana Pick Kiwifruit
Baseline 0 0 0 0 0 0
Octo 0 0 0 0 0 0
OpenVLA 0 0 0 0 33.3% 0
Ours 11.1 % 14.8 % 0 14.8 % 33.3% 0
@@
In domain Finetuning Generalization Table 4 presents the performance of the proposed Diffusion Transformer Policy compared to baseline methods. We observe different objects demonstrate various performances according to their attributes. The banana is the easiest object to pick up because the banana is longer, while kiwifruit is fat compared to other objects and all models achieve poor performance. The proposed Diffusion Transformer Policy effectively improves the diffusion action head according to Table 4. We find the discretized action head baseline achieves poor performance. Meanwhile, the Diffusion Transformer Policy is still able to pick up the novel object (e.g., the pink object) with a low success rate, while the baseline methods totally fail.
@@
Table 4: Comparision on Real-Franka tasks with few-shot fine-tuning. We represent the values in success rate. Block is “Pick up the green block”, Banana is “Pick up the banana”, Kiwi is “Pick up the Kiwifruit”, Tiny is unseen object task “Pick up the green tiny block”, Oval is unseen object task “Pick up the pink oval block”.
@@
Discretized Action Head Brohan et al. (2022) 19.3% 29.6% 51.9% 14.8% 0 0
Diffusion Action Head  Team et al. (2024) 34.8 % 40.7% 85.2% 25.9% 22.2% 0
@@
Table 5: Ablation on Maniskill2 about the number of history observation images and the length of the trajectory. #obs indicates the number of history observation images. #traj shows the length of trajectory, i.e., the sum length of observation and action prediction chunks. SingleYCB indicates PickSingleYCB, ClutterYCB indicates PickClutterYCB, SingleEGAD indicates PickSingleEGAD.
#obs #traj All PickCube StackCube SingleYCB ClutterYCB SingleEGAD
@@ 4.5 Ablation Study
Trajectory length. The length of action chunks has an important effect on the performance of different tasks. Table 5 demonstrates that the performance increases as the increase of trajectory length. Meanwhile, we notice the performance of more complex tasks, e.g., PickClutterYCB, significantly increases with the increase of trajectory length, while the easy task, e.g., PickCube, maintains high performance after the trajectory length is more than 4. Meanwhile, the long horizon optimization significantly facilitates the performance since long horizon optimization is able to provide the target object position and help the model understand the localization of the object. For example, task PickClutterYCB with multiple YCB objects, requires the model to understand which one is the corresponding object.
@@
Execution steps. Since the proposed model is able to predict multiple future actions, we can execute multiple steps in one inference. Here, we ablate the effect of execution steps using a model that has a trajectory length of 32 for quick evaluation in Table 6. The ablation study shows that the short execution steps are slightly better longer execution steps, i.e., the farther away from the current frame, the worse the prediction quality.
@@ 4.6 Visualized Comparison
Real Franka Arm. We also illustrate the comparison between the Diffusion Transformer Policy and diffusion action head baseline on real Franka Arm in Figure 5. We demonstrate the experimental results under few-shot finetuning setting. We find the proposed method achieves better action prediction when the Gripper is approaching the object and finally picks up the small green object successfully, while the baseline fails to pick up due to the inaccurate grasp position. Meanwhile, we observe that failures are usually caused by a tiny position bias and we can not even directly discriminate the position by eyes from the image. For those cases, we argue the diffusion transformer policy has learnt better grasp position during the pretraining stage, and thus reduce failure rate due to the wrong grasp pose, while it is difficult for the diffusion action head. We demonstrates more failure cases to analyze the challenges in the Franka Arm.
Figure 5: Visualized comparison between Diffusion Transformer Policy and Diffusion Action Head strategy on Real Franka Arm (Pick up the green block). The first raw is Diffusion Tranformer Policy, while the second raw is the baseline method with Diffusion Action Head.
5 Conclusion, Discussion and Future Work
@@
A limitation of the Diffusion Transformer Policy is that it requires multiple denoising steps during inference, which will impede the inference speed in the real application. In this paper, we focus on the modeling of complex and diverse actions. We think it is possible to improve the finetuning strategy with a few denoising steps to accelerate the inference speed. Meanwhile, the failure cases indicate it is important for the model to plan a right trajectory and grasp pose for picking object.

### v6

@@ 4.3 CALVIN
Comparisons. Table 2 presents the comparisons with previous methods on Calvin and the proposed methods. Without whistles and bells, the proposed Diffusion Tranformer Policy achieves the state-of-the-art results. Particularly, we only use RGB camera stream for observation. The superior demonstrates the effectiveness of Diffusion Transformer Policy. Meanwhile, the pretraining on Open X-Embodiment Datasets significantly facilitates the performance by 1.23, which demonstrates the transferability of Diffusion Transformer Policy. By contrast, the performance of diffusion action head is worse than Diffusion Transformer Policy by 0.45, though we load similar pretraining weights for diffusion head architecture. DiT Policy can perceive visual subtle nuances for long-horizon tasks, and scale across different environments, e.g., transferring the knowledge from the diverse real datasets to the CALVIN dataset.
Table 2: The comparisons with state-of-the-art approaches on Calvin Benchmark under success rate and average success length. MDT [57] is from issue 9 of its GitHub repo and GR-MG [31]. ‘S-RGB’ indicates Static RGB, ‘G-RGB’ indicate Gripper RGB. ‘S-RGBD’ indicate Static RGB-D, ‘G-RGBD’ indicates Gripper RGB-D, ‘P’ is the observation arm position ‘Proprio’, ‘Cam’ indicates camera parameters. ‘Pretraining’ indicates pretraining dataset or multi-modal model.
@@
MDT* [57] 61.7% 40.6% 23.8% 14.7% 8.7% 1.54
SPIL  [80] S-RGB,G-RGB 74.2% 46.3% 27.6% 14.7% 8.0% 1.71
RoboFlamingo [33] S-RGB,G-RGB 82.4% 61.9% 46.6% 33.1% 23.5% 2.47
SuSIE [3] S-RGB 87.0% 69.0% 49.0% 38.0% 26.0% 2.69
GR-1 [73] S-RGB,G-RGB,P 85.4% 71.2% 59.6% 49.7% 40.1% 3.06
3D Diffuser [26] S-RGBD,G-RGBD,P,Cam 92.2% 78.7% 63.9% 51.2% 41.2% 3.27
diffusion head w/o pretrain S-RGB 75.5% 44.8% 25.0% 15.0% 7.5% 1.68
diffusion head S-RGB 94.3% 77.5% 62.0% 48.3% 34.0% 3.16
Ours w/o pretrain S-RGB 89.5% 63.3% 39.8% 27.3% 18.5% 2.38
Ours S-RGB 94.5% 82.5% 72.8% 61.3% 50.0% 3.61
Table 3: Comparison with Diffusion Policy[12], Octo[69] and OpenVLA[28] on LIBERO[36]. Besides the results of our model, all of the other results are from [28]. The results are a little different with we reported in the paper, this is because that we find a more suitable scheduler for finetuning and make a little increasing on the performance.
Method LIBERO-SPATIAL LIBERO-OBJECT LIBERO-GOAL LIBERO-LONG Average
Diffusion Policy from scratch[12] 78.3% 92.5% 68.3% 50.5% 72.4%
Octo fine-tuned[69] 78.9% 85.7% 84.6% 51.1% 75.1%
OpenVLA fine-tuned[28] 84.9% 88.4% 79.2% 53.7% 76.5%
DiT Policy fine-tuned(Ours) 84.2% 96.3% 85.4% 63.8% 82.4%
4.4 LIBERO
LIBERO is a general benchmark focused on knowledge transfer in multitask and lifelong robot learning problems[36]. This benchmark consists of four sub-datasets, which are called LIBERO-SPATIAL, LIBERO-OBJECT, LIBERO-GOAL and LIBERO-100, respectively. These sub-datasets are used to test different abilities of the models. LIBERO-SPATIAL tests the ability of spatial relationship understanding, containing data of the same object set but different layouts. LIBERO-OBJECT tests the ability of object transferring, containing data of the same layouts but different object sets. LIBERO-GOAL tests the ability of task understanding and transferring, containing data of the same object sets and layouts but different tasks. LIBERO-100 can further divide into LIBERO-90 and LIBERO-10(Also called LIBERO-LONG), which are used for policy pretraining and long-horizon tasks testing, respectively. LIBERO-100 consists of data of diverse objects, layouts and backgrounds. In the latest version of OpenVLA[28], they compares the performance of several models on LIBERO. We follow OpenVLA and make a comparison with these models. We first pretrain our model on the Open X-Embodiment[50] then full-finetune it on LIBERO. We use the dataset that slightly modified by OpenVLA[28] in order to show a fair comparison. The results are shown in Table 3. Besides the results of our model, all of the other results are from [28]. We achieve a better performance on most of the sub-datasets of LIBERO and make a nearly 6% increasing on the average success rate. Specially, we make great progress on the LIBERO-LONG, which means a great potential for completing the long-horizon tasks of our model.
4.5 Maniskill2
Maniskill2 [18] is the next generation of the SAPIEN Maniskill benchmark [44], which is widely used to evaluate the generalized manipulation ability of the embodied models. It contains 20 different manipulation tasks families and over 4 million demonstration frames with different settings, including rigid/soft body, single/dual arm, etc. Maniskill2 also provides a fast and easy way to change the camera view and replay the trajectories. It is useful for the researchers working on generalized policy.
Setup. In our experiments, we select 5 tasks (PickCube-v0, StackCube-v0, PickSingleYCB-v0, PickClutterYCB-v0, PickSingleEGAD-v0) from Maniskill2, and then construct a camera pool with 300,000 random cameras, then sample 20 cameras from the camera pool to render a trajectory each time, and finally obtain about 40K trajectories totally. Given that the number of trajectories is large, we thus train the network on Maniskill2 from scratch. Moreover, we split the dataset into training set and validation set according to a ratio of 19:1. During the splitting, it is guaranteed that the single trajectory rendered under different camera views will appear in either training set or validation set in order to avoid data leaking. Specially, there are 74 different categories to pick and place in the task family PickSingleYCB-v0. In addition to the mentioned rules, we ensure each category can be found in both training set and validation set. After that, we sample 100 trajectories for each task family randomly from the validation set, constructing a close loop evaluation dataset with 500 trajectories in total. While training, considering the balance between different task families, we adjust the number of data pieces to the same by simply copying the trajectories from task family with fewer trajectories originally. The ability of the model is measured by the success rate of each task family while executing the close loop evaluation dataset.
Optimization Details. We optimize the network with AdamW [38] by 50,000 steps on Maniskill2 and we set the learning rate as 0.0001. The number of training timesteps $T$ is 100 in Maniskill2 and the global batch size is 1024.
Comparisons. Table 10 compares the proposed method with discretized action head [6, 7] and diffusion action head [69]. The experiments demonstrate Diffusion Transformer Policy achieves better results compared to Discretization Action Head strategy [6] under the large scale novel view maniskill2 benchmark. Meanwhile, Diffusion Transformer Policy demonstrates better performance in more complex tasks, e.g., Diffusion Transformer Policy improves diffusion action head [69] by 20% in task PickSingleYCB and by 12% task PickClutterYCB. Those experiments show that Diffusion Transformer Policy achieves better scalability in the large scale diverse datasets, and meanwhile achieves better generation in camera view generalization.
Table 4: Comparision with two baseline methods on Maniskill2 under success rate. SingleYCB indicates PickSingleYCB, ClutterYCB indicates PickClutterYCB, SingleEGAD indicates PickSingleEGAD. Disc ActionHead indicates Discretized Action Head strategy [6], while Diff ActionHead shows Diffusion Action Head [69].
Method All PickCube StackCube SingleYCB ClutterYCB SingleEGAD
Disc ActionHead 30.2% 41.0% 33.0% 22.0% 1.0% 54.0%
Diff ActionHead 58.6% 86.0% 76.0% 37.0% 24.0% 70.0%
DiT Policy(ours) 65.8% 79.0% 80.0% 62.0% 36.0% 72.0%
@@ 4.6 Real Franka Arm
We finally evaluate the proposed method by pretraining on Open X-Embodiment Datasets [50], and evaluate it in our Franka Arm environment under zero-shot generalization, 10-shot generalization, and in domain finetuning generalization. We meanwhile compare it by finetuning on Libero [36].
@@
Setup. We set up the franka on the table with a black background. Meanwhile, we use a single third-person RGB camera about 1.5 meters away from the Franka Arm. Please refer to Figure 4 for the visualized demonstration. Considering that the environment of our Franka setup is different from the scenes in the pretraining data Open X-Embodiment [68], we mainly evaluate the proposed method on out-of-the-box generation and few-shot generation. In our experiments, we evaluate each model with the same scene, and the object is placed in 9 similar positions in a 9-grid format in front of the franka arm. Meanwhile, we maintain a small variance in those positions placing the objects for evaluation.
@@
To evaluate the ability of quick finetuning with a few training samples, we set two pick and place tasks, and further collect five complex manipulation tasks, i.e., ‘Pick up the bowl and pour the balls into the box’, ‘Open the box and pick up the block within the box’, ‘pick up the banana and move it into the pen container’, ‘stack the bowls’, and ‘pick up the cup and pouring the coffee beans from the cup into the bowl’. We collect 10 samples for each task for 10-shot finetuning generalization. Figure 4 presents the scenes and tasks. The image in Figure 4 is the model input. Our real-world environment is challenging since the object (less than 20mm) is small compared to the whole scene. We also present a simple evaluation for in-domain evaluation for comparing different action strategies under in-domain generalization with 5 pick and place tasks, including ‘green block’, ‘kiwifruit’, ‘banana’, ‘the tiny green block’, ‘pink block’. Meanwhile, we collect 50 trajectories for each task in the first three tasks, while leaving the remaining two tasks (‘Pick up the tiny green block’ and ‘Pick up the pink block’) for out-of-distribution evaluation.
Figure 4: Illustration of Franka environment and exampled tasks.
@@
Finetuning details. In our experiments, we finetune the proposed method on the real Franka Arm with Lora [22] and AdamW [38] for 10,000 steps. We set the number of timesteps as 100 for DDPM [21], and batch size as 512.
10-shot Finetuning Generalization We directly take the models pretrained on Open X-Embodiment to evaluate 10-shot generalization in our environments. Here, to maintain the consistency with OpenVLA [28], we finetune the network with one observation and one step prediction. We compare the proposed method with OpenVLA [28] and Octo-base [69] models. Besides, we evaluate it on long horizon tasks. Table 5 presents the proposed Dit Policy archives consistent improvement compared to Octo and OpenVLA. Specifically, the proposed DiT policy demonstrates slight improvement on Pickup tasks, while the proposed DiT policy can achieve clear better performance on all more complex tasks. For those long-horizon tasks, we can find OpenVLA can complete the first task effectively while failing to under- stand the long-horizon task. For example, it completely misunderstands the insert operation, which we illustrate in the demo video. Differently, Octo is better at completing tasks with rotations and approaching the second tasks as il- lustrated with the Demo video. By contrast, the proposed DiT policy is more robust, and achieves clearly better perfor- mance in long-horizon tasks. We provide a visualized analysis in Appendix C.1.
Table 5: Comparision with Octo and Openvla on Real Franka Arm under 10-shot finetuning. Results are reported as success rate. 10-shot finetuning indicates we finetune the model with only 10 samples for each task. ‘Pourballs’ indicates pick up the bowl and pour the balls into the box. ‘Open & Pick’ indicates ‘open the box (Step-1) and pick up the block(Step-2) from the box’. ‘Pick&Insert’ indicates ‘pick up the banana (Step-1) and insert into the small pen container (Step-2)’. ‘Stacking Bowls’ is ‘stack the three bowls’, in which ‘Step-1’ is the first tacking and ‘Step-2’ is the second tacking. ‘Pick& Pouring’ is picking the mark cup (Step-1) and pouring the coffee beans from the cup into the bowl (Step-2).
Method PickBlock PickBanana Pourballs Open & Pick Stacking Bowls Pick & Insert Pick & Pouring
Step-1 Step-2 Step-1 Step-2 Step-1 Step-2 Step-1 Step-2
Octo [69] 7.4% 22.2% 14.8% 22.2% 0 10% 0 20% 0 10% 0
OpenVLA [28] 7.4% 33.3% 18.5% 7.4% 0 40% 0 80% 0 20% 0
Ours 14.8% 33.3% 29.6% 29.6% 7.4% 50% 12% 90% 10% 40% 10%
@@
In domain Finetuning Generalization Table 6 presents the performance of the proposed DiT Policy compared to baseline methods. We observe different objects demonstrate various performances according to their attributes. The banana is the easiest object to pick up because the banana is longer, while kiwifruit is fat compared to other objects and all models achieve poor performance. The proposed DiT Policy effectively improves the diffusion action head according to Table 6. We find the discretized action head baseline achieves poor performance. Meanwhile, the DiT Policy is still able to pick up the novel object (e.g., the pink object) with a low success rate, while the baseline methods totally fail.
@@
Table 6: Comparision on Real-Franka tasks with few-shot fine-tuning about different action strategies. Discretized indicates discretized action head and Diffusion indicates Diffusion action head. We represent the values in success rate. Block is ‘Pick up the green block’, Banana is ‘Pick up the banana’, Kiwi is ‘Pick up the Kiwifruit’, Tiny is unseen object task ‘Pick up the green tiny block’, Oval is unseen object task ‘Pick up the pink oval block’.
@@
Discretized [6] 19.3% 29.6% 51.9% 14.8% 0 0
Diffusion  [69] 34.8 % 40.7% 85.2% 25.9% 22.2% 0
@@
Table 7: Ablation on Maniskill2 about the number of history observation images and the length of the trajectory. #obs indicates the number of history observation images. #traj shows the length of trajectory, i.e., the sum length of observation and action prediction chunks. PickC indicates PickCube, StackC indicates StackCube, SingleYCB indicates PickSingleYCB, ClutterYCB indicates PickClutterYCB, EGAD indicates PickSingleEGAD.
#obs #traj All PickC StackC SingleYCB ClutterYCB EGAD
@@ 4.7 Ablation Study
Trajectory length. The length of action chunks has an important effect on the performance of different tasks. Table 7 shows that performance increases with increasing trajectory length. Meanwhile, we notice the performance of more complex tasks, e.g., PickClutterYCB, increases significantly with increasing trajectory length, while the easy task, e.g., PickCube, maintains high performance after the trajectory length is greater than 4. Meanwhile, the long horizon optimization significantly facilitates the performance since long horizon optimization is able to provide the target object position and help the model understand the localization of the object. For example, task PickClutterYCB with multiple YCB objects, requires the model to understand which one is the corresponding object.
@@
Execution steps. Since the proposed model is able to predict multiple future actions, we can execute multiple steps in one inference. Here, we ablate the effect of execution steps under a model with trajectory length 32 in Table 8. The ablation study shows that the short execution steps are slightly better longer execution steps, i.e., the farther away from the current frame, the worse the prediction quality.
@@ 4.8 Visualized Comparison
Real Franka Arm. We also illustrate the comparison between the Diffusion Transformer Policy and diffusion action head baseline on real Franka Arm in Figure 10. We demonstrate the experimental results under few-shot finetuning setting. We find the proposed method achieves better action prediction when the Gripper is approaching the object and finally picks up the small green object successfully, while the baseline fails to pick up due to the inaccurate grasp position. Meanwhile, we observe that failures are usually caused by a tiny position bias and we can not even directly discriminate the position by eyes from the image. For those cases, we argue the diffusion transformer policy has learnt better grasp position during the pretraining stage, and thus reduce failure rate due to the wrong grasp pose, while it is difficult for the diffusion action head. We demonstrates more comparisons to analyze the challenges in the Franka Arm.
Figure 6: Visualized comparison between Diffusion Transformer Policy and Diffusion Action Head strategy on Real Franka Arm (Pick up the green block). The first raw is DiT Policy, while the second raw is the baseline method with Diffusion Action Head.
5 Conclusion and Discussion
@@
A limitation of the Diffusion Transformer Policy is that it requires multiple denoising steps during inference, which will impede the inference speed in the real application. In this paper, we focus on the modeling of complex and diverse actions. We think it is possible to improve the finetuning strategy with a few denoising steps to accelerate the inference speed.
Supplementary Material